OpenStudy (anonymous):

How do I evaluate: 3^-3 + 3^-3/3^-5?

OpenStudy (anonymous):

\[\frac{1}{3^3}+3^2\] is a start

OpenStudy (anonymous):

Recall that \[x^{-n}=\dfrac{1}{x^n}\]Thus, we have:\[\dfrac{1}{3\cdot3\cdot3}+\dfrac{\left(\dfrac{1}{3\cdot3\cdot3}\right)}{\left(\dfrac{1}{3\cdot3\cdot3\cdot3\cdot3}\right)}\]\[\dfrac{1}{3\cdot3\cdot3}+\dfrac{3\cdot3\cdot3\cdot3\cdot3}{3\cdot3\cdot3}\]\[\dfrac{3\cdot3\cdot3\cdot3\cdot3 + 1}{3\cdot3\cdot3}\]\[\dfrac{3^5+1}{3^3}\]